# Significant impact of Galactic dark matter particles on annihilation signals from Sagittarius analogues

### v1

@@ Abstract
We examine the gamma-ray signal from dark matter (DM) annihilation from analogues of the Sagittarius (Sgr) dwarf spheroidal galaxy in the Auriga cosmological simulations. For velocity-dependent annihilation cross sections, we compute emissions from simulated Sgr subhalos and from the Milky Way (MW) foreground. In addition to the annihilation signals from DM particles bound to Sgr, we consider for the first time the annihilation of DM particles bound to the MW that overlap spatially with Sgr. For p-wave models this contribution can enhance the signal by over an order of magnitude, while for d-wave models the enhancement can be over two orders of magnitude. For Sommerfeld models, the corresponding emission decreases by up to nearly an order of magnitude. For Sommerfeld and s-wave models, the Sgr source can be visible above the MW foreground emission, while for p and d-wave models, the signal towards Sgr is most likely dominated by foreground MW emission. We interpret our results within the context of the observed gamma-ray emission from Sgr, finding that the templates from simulations likely have spatial morphology that is too extended to explain the point-like emission that is observed.
@@ 1 Introduction
In this paper, we examine predictions for the DM annihilation signal from Sgr using analogues in the Auriga cosmological simulations. We provide predictions for the signals for both velocity-independent and velocity-dependent models. We consider, for the first time, the contribution to the DM annihilation signal from DM particles bound to the MW potential that spatially overlay with the Sgr analogues. For the purposes of this paper, we refer to this latter component as “unbound” particles, since they are not bound to the Sgr analogues. We demonstrate that the contribution from these unbound particles has a significant impact on the DM annihilation signals for the velocity-dependent models. We quantify the range of enhancement correction factors from this component and examine the implications for the sensitivity to velocity-dependent models using Fermi-LAT data.
@@ 2.1 The Auriga Simulations
We use the high resolution level 3 Auriga simulations in our study. At this resolution level, the DM particle mass is $m_{\rm DM}\sim 4\times 10^{4}~{}{\rm M}_{\odot}$, the baryonic particle mass is $m_{b}=5\times 10^{3}~{}{\rm M}_{\odot}$, and the Plummer equivalent gravitational softening length is $\epsilon=184$ pc [31, 32]. The six MW analogues at the level 3 simulations studied in this work are Au6, Au16, Au21, Au23, Au24 and Au27.
@@ 2.2 Identifying Sagittarius analogues
Host name Dwarf $D$ [kpc] $\log_{10}$($M_{*}$${}_{\odot}$) $\log_{10}$($M_{\rm tot}$${}_{\odot}$) $R_{\text{max}}$ [kpc] Angular size [deg] Au16 9 28.08 7.933 8.810 1.957 3.945 Au21 10 44.70 8.446 8.780 1.078 1.395 Au23 4 49.01 8.297 8.851 1.583 1.884 Au23 7 38.11 8.047 8.790 1.654 2.502 Au24 24 60.13 7.503 8.629 2.970 2.788 Au27 25 35.91 7.598 8.170 0.852 1.390
@@
Table 1: Summary of the properties of the six simulated Sgr-like systems in Auriga. From left to right, the columns list the Auriga halo ID of the MW analogue hosts, the dwarf ID of the Sgr analogues, the distance of the observer to the dwarf, the bound stellar mass, total bound mass, radius $R_{\rm max}$, and the angular size of the Sgr analogues. The angular size is calculated using the distance from the halo to the sun.
@@
Table 1 lists the properties of the six selected Sgr analogues. The columns of the table from left to right list the Auriga halo ID of the MW analogue hosts, the dwarf ID of the Sgr analogues, distance of the Sgr analogues to the observer, the bound stellar mass of the Sgr analogue, the total (DM and stars) bound mass of the Sgr analogue, the radius, $R_{\rm max}$, at which the circular velocity of the Sgr analogue reaches its maximum value, and the angular size of the Sgr analogues as seen from the Solar position.
@@
While the definition of Sgr analogues in ref. [16] relies on the stellar mass bound to the progenitor at present day, it can be useful to also consider other properties like morphology of tidal debris and orbital trajectory. Figure 1 shows the distribution of stellar mass for each Sgr analogue, including mass that has been lost after accreting onto the MW host. It is clear that the selection criteria results in a variety of possible morphologies, largely driven by the total mass at infall and the orbit of the analogues. Sgr has an estimated total stellar mass (prior to infall) of $M_{\ast}\sim 10^{8.3}~{}{\rm M}_{\odot}$ [33] and eccentricity $e\equiv(r_{\text{apo}}-r_{\text{peri}})/(r_{\text{apo}}+r_{\text{peri}})\sim 0.65$ [34]. Of the analogues listed in table 1, Sgr 9 and Sgr 7 fall into a reasonable range of matches ($10^{8.1}~{}{\rm M}_{\odot}<M_{\ast}<10^{8.6}~{}{\rm M}_{\odot}$ and $0.750.6<e<$ 0.75). An analysis of the full population of stellar streams in the Auriga simulations is forthcoming [35, 36].
@@
Figure 1: Stellar mass maps of the tidal debris associated with each Sgr analogue identified in Auriga. From left to right are Sgr(9, 10, 4, 7, 24, 25) as denoted in table 1, with each 500 kpc per side panel centred on the MW host. As extreme examples, Sgr 10 is incredibly massive ($M_{\ast}\sim 10^{9.5}$ ${}_{\odot}$) and on a very radial orbit $(e\sim 0.9)$, while Sgr 4 is on a nearly circular orbit (with eccentricity $e\sim 0.1$). In terms of stellar mass and orbital trajectory, Sgr 9 and Sgr 7 are the closest matches to the real Sgr (see Section 2.2 for further discussion).
@@ 2.3 Finding the Solar positions
For each MW-Sgr system, we choose the Solar position which places the Galactic center ($l=0^{\circ}$, $b=0^{\circ}$) and the Sgr analogue ($l=5.569^{\circ}$, $b=-14.166^{\circ}$) at the correct on-sky locations. To determine the Solar position, we sample over spherical shells of 0.01 kpc radial width that are at a distance ranging from $7-9$ kpc from the galactic center. Within each spherical shell, the position is checked for 500 points by varying the azimuthal angle between 0 and $2\pi$, and 500 points by varying the inclination angle between 0 to $\pi$. For each system, we find a Solar position that places the Sgr analogue within one degree of the targeted latitude and longitude. The resulting distances between the best Solar position and its associated Sgr analogue are given in the third column of table 1.
@@ 4.1 Including both bound and unbound particles
The $\mathcal{J}$-factors are computed separately for the smooth halo and subhalos. For the smooth halo component, the local DM density around each DM particle is estimated using a Voronoi tessellation. Then the DM relative velocity distribution is computed at each point on a spherical mesh using the nearest 500 DM particles [24]. These values are then interpolated to obtain an estimate for the DM relative velocity distribution at every point in the smooth halo. The $\mathcal{J}$-factors are then computed by integrating the contribution from all DM particles along the line-of-sight to the Sgr analogue.
For subhalos (including the Sgr analogues), the local DM density is found using the best fit Einasto density profile within $R_{\rm max}$ and a Voronoi tessellation outside of $R_{\rm max}$. The DM relative velocity distribution is calculated using a Maxwell-Boltzmann (MB) distribution given by
$P_{\rm MB}(|{\bf v}_{\rm rel}|)=\sqrt{\frac{2}{\pi}}\frac{v_{\rm rel}^{2}}{\sigma^{3}}\exp{\left(-\frac{v_{\rm rel}^{2}}{2\sigma^{2}}\right)},$ (4.4)
where the 1D relative velocity dispersion, $\sigma$, is set equal to the mean relative velocity dispersion of the subhalo. The mean relative velocity dispersion is found by calculating the DM relative velocity distribution in spherical shells, fitting each of these with a MB distribution and finding the mean velocity dispersion over all shells. The mean velocity dispersion is independent of position within the subhalo. The resulting MB distribution is then used to compute the effective $\mathcal{J}$-factors using eq (4.2).
@@
Figure 4 shows the $\mathcal{J}$-factor maps for a region around the Sgr 9 analogue, for the different velocity-dependent DM annihilation models. For each panel we use a pixel size of 0.3 degree on each side. The top panels show the $\mathcal{J}$-factor maps for only the Sgr 9 analogue including bound and unbound particles, while the bottom panels show the $\mathcal{J}$-factor maps for Sgr 9 once the contribution from the MW smooth halo and other subhalos in the region are added in. We choose to highlight Sgr 9 as it is the Sgr analogue with the largest enhancement factors for the p and d-wave models, and the largest decrement in emission for the Sommerfeld model, when we account for the unbound particles (see table 2 below). As shown in the figure, the MW diffuse signal from the smooth halo dominates the source signal in all cases but the Sommerfeld case, where the velocity dependence goes as $1/v$. For the Sommerfeld case, we see that the source is clearly visible above the MW emission. In the s-wave case, the source is much less prominent, and for the p and d-wave cases, the source is not visible above the MW emission.
Figure 4: $\mathcal{J}$-factor maps for a region of interest centered on the Sgr 9 analogue for the Sommerfeld, s-wave, p-wave, and d-wave models. The top panels show only the contribution from the Sgr 9 analogue, including both bound and unbound DM particles. The bottom panels show the contribution from the entire region, including the MW foreground emission. $\alpha$ and $\beta$ are the polar and azimuthal angles in a randomly defined spherical coordinate system centered on the Sgr analogue, with the $z$-axis pointing from the Sun’s position to the Sgr analogue. The dashed white circle specifies the angular extension of the Sgr source. The bright component in the upper right corner of each map in the bottom panels is the emission from DM annihilation from the Galactic center. Only in the Sommerfeld case is the emission from the Sgr source clearly visible above the diffuse emission from the MW.
@@ 4.2 Including only bound particles
It is interesting to compare the true $\mathcal{J}$-factors computed from the bound and unbound DM particles in the Sgr analogues to the hypothetical case in which we only consider the emission from DM particles that are classified as bound to the Sgr analogues. For this comparison, we exclude the particles from the $\mathcal{J}$-factor calculation that spatially overlap with the Sgr analogues, but are bound to the MW host halos and not to the Sgr analogues. We then compare the results to the $\mathcal{J}$-factors calculated using the full simulation. The spherically-averaged $\mathcal{J}$-factors as a function of angle from the center of the Sgr analogues are shown in the right panels of figure 3, where only the bound particles to the Sgr analogues are included. Comparing to the left panel, the largest relative enhancements due to the unbound DM particles in the Sgr region come from the d-wave component, although the total emission is still well below that of the MW emission in this case. It is interesting to note that for the case of p and d-waves, there is a larger scatter when including the unbound DM particles. For the Sommerfeld case, there is a reduction relative to the foreground when all DM particles are included, though the scatter is similar in both cases. For s-wave, including the unbound particles has minimal effect, as expected.
@@
Au16 9 0.333 1.186 15.015 190.072
Au21 10 0.826 1.008 1.504 2.243
Au23 4 0.686 1.006 2.164 4.655
Au23 7 0.570 1.019 3.260 10.427
Au24 24 0.335 1.104 12.001 130.466
Au27 25 0.348 1.048 9.510 86.337
Table 2: Modifications to the $\mathcal{J}$-factors of the Sgr analogues due to the unbound DM particles. Each entry shows the ratio of the $\mathcal{J}$-factor for the full simulation results (including both the bound and unbound DM particles) to the case in which only the DM particles bounds to the Sgr analogues are considered. Note that there is an enhancement for p and d-waves, while there is a reduction for the Sommerfeld case. There is essentially no change in the s-wave case.
The enhancement factors when including the unbound DM particle in the calculation of the $\mathcal{J}$-factors of the Sgr analogues for the p and d-waves, and the corresponding reduction in the $\mathcal{J}$-factors for the Sommerfeld case are shown in table 2. Including the unbound DM particles can increase the $\mathcal{J}$-factors of the Sgr analogues by over an order of magnitude for the p-wave model, and by over than two orders of magnitude for the d-wave model. For the Sommerfeld model, the $\mathcal{J}$-factors decrease by up to nearly an order of magnitude, while for the s-wave model, the $\mathcal{J}$-factors minimally change.
Figure 5: $\mathcal{J}$-factor maps for a region of interest centered on the Sgr 7, Sgr 9, Sgr 24, and Sgr 25 analogues for the Sommerfeld model, including the MW foreground emission. In the top panels, both the bound and unbound DM particles are included for the Sgr analogues, while in the bottom panels only the bound DM particles are included for the Sgr analogues.
Figure 5 shows a comparison of the $\mathcal{J}$-factor maps for four Sgr analogues for the Sommerfeld model including both the bound and unbound DM particles (top panels) and only the DM particles bound to the Sgr analogues (bottom panels). Emission from the MW foreground and from all other subhalos in the line of sight is also shown in the maps. The four Sgr analogues shown in the figure have the largest suppression of Sommerfeld annihilation due to the inclusion of unbound DM particles. Comparing the top and bottom panels of the figure, we can clearly see that the source is fainter when both the bound and unbound DM particles are included in the maps.

### v2

@@ Abstract
We examine the gamma-ray signal from dark matter (DM) annihilation from analogues of the Sagittarius (Sgr) dwarf spheroidal galaxy in the Auriga cosmological simulations. For velocity-dependent annihilation cross sections, we compute emissions from simulated Sgr subhalos and from the Milky Way (MW) foreground. In addition to the annihilation signals from DM particles bound to Sgr, we consider for the first time the annihilation of DM particles bound to the MW that overlap spatially with Sgr. For p-wave models this contribution can enhance the signal by over an order of magnitude, while for d-wave models the enhancement can be over three orders of magnitude. For Sommerfeld and s-wave models, the corresponding emission does not significantly change. For the Sommerfeld model, the Sgr source can be visible above the MW foreground emission, while for s, p and d-wave models, the signal towards Sgr is most likely dominated by foreground MW emission. We interpret our results within the context of the observed gamma-ray emission from Sgr. We find that, given the background emission estimated from this region, the templates from simulations likely have spatial morphology that is too extended to explain the point-like emission that is observed.
@@ 1 Introduction
In this paper, we examine predictions for the DM annihilation signal from Sgr using analogues in the Auriga cosmological simulations. We provide predictions for the signals for both velocity-independent and velocity-dependent models. We consider, for the first time, the contribution to the DM annihilation signal from DM particles bound to the MW potential that spatially overlay with the Sgr analogues. For the purposes of this paper, we refer to this latter component as “unbound” particles, since they are not bound to the Sgr analogues. We demonstrate that the contribution from these unbound particles has a significant impact on the DM annihilation signals for the velocity-dependent models. We quantify the range of enhancement factors from this component and examine the implications for the sensitivity to velocity-dependent models using Fermi-LAT data.
@@ 2.1 The Auriga Simulations
We use the high resolution level 3 Auriga simulations in our study. At this resolution level, the DM particle mass is $m_{\rm DM}\sim 5\times 10^{4}~{}{\rm M}_{\odot}$, the baryonic particle mass is $m_{b}=5\times 10^{3}~{}{\rm M}_{\odot}$, and the Plummer equivalent gravitational softening length is $\epsilon=184$ pc [31, 32]. The six MW analogues at the level 3 simulations studied in this work are Au6, Au16, Au21, Au23, Au24 and Au27.
@@ 2.2 Identifying Sagittarius analogues
Host Dwarf $D$ [kpc] $\log_{10}$($M_{*}$/M) $\log_{10}$($M_{\rm tot}$/M) $R_{\text{max}}$ [kpc] Angular size [deg] eccentricity Au16 9 28.08 7.933 8.810 1.957 3.945 0.65 Au21 10 44.70 8.446 8.780 1.078 1.395 0.90 Au23 4 49.01 8.297 8.851 1.583 1.884 0.08 Au23 7 38.11 8.047 8.790 1.654 2.502 0.67 Au24 24 60.13 7.503 8.629 2.970 2.788 0.62 Au27 25 35.91 7.598 8.170 0.852 1.390 0.66 Sgr $26.4$ $8.6$ $\gtrsim 1$ $\gtrsim 12.2$ 0.65
@@
Table 1: Summary of the properties of the six simulated Sgr-like systems in Auriga. From left to right, the columns list the Auriga halo ID of the MW analogue hosts, the dwarf ID of the Sgr analogues, the distance of the observer to the dwarf, the present-day bound stellar mass, total present-day bound mass, radius $R_{\rm max}$, the angular size, and the eccentricity of the Sgr analogues. The angular size is calculated using the distance from the halo center to the Sun. The bottom row gives the properties of the observed Sgr dwarf.
@@
Table 1 lists the properties of the six selected Sgr analogues. The columns of the table from left to right list the Auriga halo ID of the MW analogue hosts, the dwarf ID of the Sgr analogues, distance of the Sgr analogues to the observer, the bound stellar mass of the Sgr analogue, the total (DM and stars) bound mass of the Sgr analogue, the radius, $R_{\rm max}$, at which the circular velocity of the Sgr analogue reaches its maximum value, the angular size of the Sgr analogues as seen from the Solar position, and the eccentricity of the Sgr analogues. In the bottom row we present the properties of the observed Sgr dwarf.
@@
While the definition of Sgr analogues in ref. [16] relies on the stellar mass bound to the progenitor at present day, it can be useful to also consider other properties like morphology of tidal debris and orbital trajectory. Figure 1 shows the distribution of stellar mass for each Sgr analogue, including mass that has been lost after accreting onto the MW host. It is clear that the selection criteria results in a variety of possible morphologies, largely driven by the total mass at infall and the orbit of the analogues. Sgr has an estimated total stellar mass (prior to infall) of $M_{\ast}\sim 10^{8.3}~{}{\rm M}_{\odot}$ [33] and eccentricity $e\equiv(r_{\text{apo}}-r_{\text{peri}})/(r_{\text{apo}}+r_{\text{peri}})\sim 0.65$ [34]. Of the analogues listed in table 1, Sgr 9 and Sgr 7 fall into a reasonable range of matches, with stellar masses prior to infall of $10^{8.5}~{}{\rm M}_{\odot}<M_{\ast}<10^{8.7}~{}{\rm M}_{\odot}$ and eccentricities $0.750.6<e<$ 0.75. Note that these infall stellar masses are different from the present-day stellar masses quoted in table 1. An analysis of the full population of stellar streams in the Auriga simulations is forthcoming [35, 36].
@@
Figure 1: Stellar mass maps of the tidal debris associated with each Sgr analogue identified in Auriga. From left to right are Sgr(9, 10, 4, 7, 24, 25) as denoted in table 1, with each 500 kpc per side panel centred on the MW host. As extreme examples, Sgr 10 is incredibly massive ($M_{\ast}\sim 10^{9.5}$ M) and on a very radial orbit $(e\sim 0.9)$, while Sgr 4 is on a nearly circular orbit (with eccentricity $e\sim 0.1$). In terms of stellar mass and orbital trajectory, Sgr 9 and Sgr 7 are the closest matches to the real Sgr (see Section 2.2 for further discussion).
@@ 2.3 Finding the Solar positions
In the simulations, the Solar position is not determined a priori. For each MW-Sgr system, we choose the Solar position such that the position of the Sgr analogue matches its correct on-sky location ($l=5.569^{\circ}$, $b=-14.166^{\circ}$), with the galactic center at $l=0^{\circ}$, $b=0^{\circ}$. To determine the Solar position, we use a temporary spherical coordinate system centered on the galactic center and sample over spherical shells of 0.01 kpc radial width that are at a galactocentric distance ranging from $7-9$ kpc. Within each spherical shell centered on the galactic center, the position is checked for 500 points by varying the azimuthal angle between 0 and $2\pi$, and 250 points by varying the inclination angle between 0 to $\pi$. Since there is no well-defined disk for the host halo galaxy, we do not enforce a constraint to keep the Sun in the galactic disk. We only constrain its distance to the galactic center. For each system, we find the Sgr analogue positions are within 0.2 degree of the targeted latitude and longitude. The resulting distances between the best Solar position and its associated Sgr analogue are given in the third column of table 1. We find that the range of circular velocities, $v_{c}=\sqrt{GM(<r)/r}$ at the Solar position for each host analogue is in the range $[233-270]$ km/s, which is consistent with the observed value [37]. The mass, $M(<r)$, is computed as the sum of the masses of dark matter, gas, and stars within radius $r$ from the center of the host halos.
@@ 4.1 Including both bound and unbound particles
The $\mathcal{J}$-factors are computed separately for the smooth halo and subhalos. For the smooth halo component, the local DM density around each DM particle is estimated using a Voronoi tessellation. Then the DM relative velocity distribution is computed at each point on a spherical mesh using the nearest 500 DM particles [24]. These values are then interpolated to obtain an estimate for the DM relative velocity distribution at every point in the smooth halo. The $\mathcal{J}$-factors are then computed by integrating the contribution from all DM particles along the line-of-sight to the Sgr analogue. For subhalos (including the Sgr analogues), the local DM density is found using the best fit Einasto density profile within $R_{\rm max}$ and a Voronoi tessellation outside of $R_{\rm max}$. The density profile is computed such that each shell has 100 DM particles. The DM relative velocity distribution is also calculated in non-uniform consecutive spherical shells from the center of the subhalo, where each shell contains 100 DM particles. This allows us to capture the radial dependence of the DM relative velocities in the subhalo. The resulting velocity distribution is then used to compute the effective $\mathcal{J}$-factors using eq (4.2).
@@
An important quantity for DM indirect detection is the $\mathcal{J}$-factor integrated over solid angle, which is given by
$\widetilde{\mathcal{J}}(\theta)=2\pi\int_{0}^{\theta}\mathcal{J}(\theta^{\prime})\sin\theta^{\prime}d\theta^{\prime}.$ (4.4)
In the left panel of figure 4, we show the integrated $\mathcal{J}$-factor profiles for the Sgr analogues including both bound and unbound DM particles. Same as in figure 3, the thin lines show the results for different Sgr analogue, while the thick lines indicate the average across all Sgr analogues. The shaded bands show the standard deviation around the mean. The red, blue, orange, and green lines correspond to the Sommerfeld, s-wave, p-wave, and d-wave models, respectively. The integrated $\mathcal{J}$-factors for the s-wave reach values at large $\theta$ which are consistent with the $\mathcal{J}$-factor computed for the observed Sgr [13].
Figure 5 shows the $\mathcal{J}$-factor maps for a region around the Sgr 9 analogue, for the different velocity-dependent DM annihilation models. For each panel we use a pixel size of 0.3 degree on each side. The top panels show the $\mathcal{J}$-factor maps for only the Sgr 9 analogue including bound and unbound particles, while the bottom panels show the $\mathcal{J}$-factor maps for Sgr 9 once the contribution from the MW smooth halo and other subhalos in the region are added in. We choose to highlight Sgr 9 as it is the Sgr analogue with the largest angular size, and is therefore most easily observed in the maps. As shown in the figure, the MW diffuse signal from the smooth halo dominates the source signal in all cases but the Sommerfeld case, where the velocity dependence goes as $1/v$. For the Sommerfeld case, we see that the source is clearly visible above the MW emission. For the s, p and d-wave cases, the source is not visible above the MW emission.
Figure 4: Integrated $\mathcal{J}$-factor profiles (eq. (4.4)) for the Sgr analogues. In the left panel the $\mathcal{J}$-factor are calculated with the bound and unbound DM particles, while in the right panel only the bound DM particles were included in the calculation. In both panels, the thin coloured lines show the results for each Sgr analogue, while the thick coloured lines correspond to the average across all Sgr analogues. The shaded bands indicate the standard error around the mean. The red, blue, orange, and green coloured lines show the results for the Sommerfeld, s-wave, p-wave, and d-wave DM annihilation models, respectively.
Figure 5: $\mathcal{J}$-factor maps for a region of interest centered on the Sgr 9 analogue for the Sommerfeld, s-wave, p-wave, and d-wave models. The top panels show only the contribution from the Sgr 9 analogue, including both bound and unbound DM particles. The bottom panels show the contribution from the entire region, including the MW foreground emission. $\alpha$ and $\beta$ are the polar and azimuthal angles in a randomly defined spherical coordinate system centered on the Sgr analogue, with the $z$-axis pointing from the Sun’s position to the Sgr analogue. The dashed white circle specifies the angular extension of the $R_{\rm max}$ of the Sgr source. The bright component in the upper right corner of each map in the bottom panels is the emission from DM annihilation from the Galactic center. Only in the Sommerfeld case is the emission from the Sgr source clearly visible above the diffuse emission from the MW. In the bottom panels, the black contours are placed at fractions of the maximum pixel value given by 0.02, 0.005 and 0.0015.
@@ 4.2 Including only bound particles
It is interesting to compare the true $\mathcal{J}$-factors computed from the bound and unbound DM particles in the Sgr analogues to the hypothetical case in which we only consider the emission from DM particles that are classified as bound to the Sgr analogues. For this comparison, we exclude the particles from the $\mathcal{J}$-factor calculation that spatially overlap with the Sgr analogues, but are bound to the MW host halos and not to the Sgr analogues. We then compare the results to the $\mathcal{J}$-factors calculated using the full simulation. The spherically-averaged $\mathcal{J}$-factors as a function of angle from the center of the Sgr analogues are shown in the right panels of figure 3, where only the bound particles to the Sgr analogues are included. Comparing to the left panel, the largest relative enhancements due to the unbound DM particles in the Sgr region come from the d-wave component, although the total emission is still well below that of the MW emission in this case. It is interesting to note that for the case of p and d-waves, there is a larger scatter when including the unbound DM particles. For the Sommerfeld case, the effect of the high speed unbound DM particles is suppressed due to the $1/v$ dependence of this model, and thus they have minimal effect on the $\mathcal{J}$-factors. The effect is also minimal for s-wave since there is no velocity dependence and the density of unbound DM particles is low. Similar results are found for the integrated $\mathcal{J}$-factors when comparing the left panel of figure 4 to the right panel, which only considers the bound DM particles in the calculation.
@@
Au16 9 1.28 1.22 16.79 1379.01
Au21 10 1.06 1.03 1.68 15.76
Au23 4 1.09 1.05 2.85 103.19
Au23 7 1.15 1.07 5.15 342.22
Au24 24 1.49 1.50 29.83 3878.96
Au27 25 1.24 1.11 15.12 2412.44
Table 2: Modifications to the $\mathcal{J}$-factors of the Sgr analogues due to the unbound DM particles. Each entry shows the ratio of the $\mathcal{J}$-factor for the full simulation results (including both the bound and unbound DM particles) to the case in which only the DM particles bounds to the Sgr analogues are considered. The ratios specify the average enhancements over all pixels incorporating each Sgr analogue. Note that there is a considerable enhancement for p and d-waves, while there is essentially no change in the Sommerfeld and s-wave cases.
Figure 6: Ratio of the $\mathcal{J}$-factors of the Sgr analogues calculated with both bound and unbound DM particles to those calculated with only the bound DM particles. In the left panel, the ratios are shown for the differential $\mathcal{J}$-factors, while in the right panel they are shown for the integrated $\mathcal{J}$-factors. In both panels, the thin coloured lines show the results for each Sgr analogue, while the thick coloured lines correspond to the average across all Sgr analogues, with the shaded band indicating the standard error around the mean. The red, blue, orange, and green coloured lines show the results for the Sommerfeld, s-wave, p-wave, and d-wave models, respectively.
The enhancement factors when including the unbound DM particles in the calculation of the $\mathcal{J}$-factors of the Sgr analogues for the four DM annihilation models are shown in table 2. The values shown are the average enhancement factors over all pixels that make up each respective Sgr analogue. Including the unbound DM particles can increase the $\mathcal{J}$-factors of the Sgr analogues by over an order of magnitude for the p-wave model, and by over three orders of magnitude for the d-wave model. For the Sommerfeld and s-wave models, the $\mathcal{J}$-factors minimally change.
The angular dependence of the enhancement factors can be seen in figure 6, which shows the ratio of the $\mathcal{J}$-factors calculated with the bound and unbound DM particles to those calculated with only the bound DM particles. We define $\theta_{\rm max}$ as the angular extent of the subhalo at radius $R_{\rm max}$. The differential $\mathcal{J}$-factor profile ratios are shown in the left panel and the integrated $\mathcal{J}$-factor profile ratios are shown in the right panel. The line colors, styles, and shaded bands are the same as in figures 3 and 4. While the enhancement factors are significant for p and d-wave models for all angles and for both the differential and integrated $\mathcal{J}$-factors, they remain negligible for the Sommerfeld and s-wave models. The ratios of the differential $\mathcal{J}$-factors in the left panel peak slightly below $R_{\rm max}$, due to the decreasing contributions from the lines-of-sight towards the edge of the subhalo.